# Tora: Trajectory-oriented Diffusion Transformer for Video Generation

### v3

@@ Abstract
Recent advancements in Diffusion Transformer (DiT) have demonstrated remarkable proficiency in producing high-quality video content. Nonetheless, the potential of transformer-based diffusion models for effectively generating videos with controllable motion remains an area of limited exploration. This paper introduces Tora, the first trajectory-oriented DiT framework that concurrently integrates textual, visual, and trajectory conditions, thereby enabling scalable video generation with effective motion guidance. Specifically, Tora consists of a Trajectory Extractor (TE), a Spatial-Temporal DiT, and a Motion-guidance Fuser (MGF). The TE encodes arbitrary trajectories into hierarchical spacetime motion patches with a 3D video compression network. The MGF integrates the motion patches into the DiT blocks to generate consistent videos that accurately follow designated trajectories. Our design aligns seamlessly with DiT’s scalability, allowing precise control of video content’s dynamics with diverse durations, aspect ratios, and resolutions. Extensive experiments demonstrate Tora’s excellence in achieving high motion fidelity, while also meticulously simulating the intricate movement of the physical world.
Introduction
Figure 1: More generated samples. Tora accommodates diverse conditions such as text-only inputs, single starting frames, and combinations of initial and final frames (as illustrated in the fifth row). It effectively manages multiple trajectories, allowing for precise manipulation of several objects. Furthermore, Tora supports video generation across different aspect ratios, resolutions, and durations, ensuring flexible content creation. For video demonstrations, please refer to our project page.
@@ Introduction
Diffusion models (Dhariwal and Nichol 2021; Ramesh et al. 2022) have demonstrated their capability to generate diverse and high-quality images or videos. Previously, video diffusion models (Ho et al. 2022b; Blattmann et al. 2023; Zhang et al. 2023a) predominantly employed U-Net architectures (Olaf Ronneberger 2015), focusing primarily on synthesizing videos of limited duration, typically around two seconds, and were constrained to fixed resolutions and aspect ratios. Recently, Sora (Brooks et al. 2024), a text-to-video generation model leveraging Diffusion Transformer (DiT) (Peebles and Xie 2023), has showcased video generation capabilities that significantly outstrip current state-of-the-art methods. Sora excels not only in the production of high-quality videos ranging from 10 to 60 seconds, but also distinguishes itself through its capacity to handle diverse resolutions, various aspect ratios, adherence to the laws of actual physics.
@@
Video generation requires consistent motion across image sequences, underscoring the importance of motion control. Previous works, such as VideoComposer (Wang et al. 2023) and DragNUWA (Yin et al. 2023), have implemented generalized motion manipulation through motion vectors and trajectories. Building on this foundation, MotionCtrl (Wang et al. 2024b) innovates by independently managing camera and object motions, thereby expanding the diversity of achievable motion patterns. Despite their promising controllable motion quality, U-Net methods are restricted to generating videos of only 16 frames at a fixed, lower resolution. This limitation hinders the smooth portrayal of motion, particularly during significant positional shifts in the provided trajectory, leading to distortion and unnatural movements, such as parallel drifting, which diverge from real-world dynamics. Consequently, there is an urgent need for a model capable of producing longer videos with robust motion control and detailed physical representations.
@@
To address these challenges, we present Tora, the first DiT model that simultaneously integrates text, images, and trajectories, enabling scalable video generation with robust motion control. Notably, our work adopts OpenSora (Zheng et al. 2024), an open-source version of Sora, as the foundational DiT model. To align motion control with the scalability of the DiT framework, we propose two novel modules: the Trajectory Extractor (TE), which converts arbitrary trajectories into hierarchical spacetime motion patches, and the Motion-guidance Fuser (MGF), designed to seamlessly integrate these patches within the stacked DiT blocks. More specifically, TE initially converts positional displacements along trajectory into the RGB domain via flow visualization techniques. These visualized displacements undergo Gaussian filtering to mitigate scattered issues. Subsequently, a 3D Variational Autoencoder (VAE) (Kingma and Welling 2013) encodes trajectory visualizations into spacetime motion latents, which share the same latent space with video patches. The motion latents are then decomposed into multiple levels of motion conditions via stacked lightweight modules. Our VAE architecture is inspired by MAGVIT-v2 (Yu et al. 2023b) but simplified by foregoing codebook dependencies. The MGF integrates adaptive normalization layers (Perez et al. 2018) to infuse multi-level motion conditions into the corresponding DiT blocks. We explored various adaptations of transformer blocks including adaptive layer normalization, cross-attention, and extra channel connections to inject the motion conditions. Among these, adaptive layer normalization emerged as the most effective to generate consistent videos following trajectory.
@@
During training, we adapt OpenSora’s workflow to generate high-quality video-text pairs and utilize an optical flow estimator (Xu et al. 2023) for trajectory extraction. We also integrate a motion segmentor (Zhao et al. 2022) with a camera detector [^1] https://github.com/antiboredom/camera-motion-detector to filter out instances dominated by camera motion, thereby improving our tracking of specific object trajectories. This careful selection process results in a dataset of 630k high-quality videos with consistent motion. With an adapter-like strategy (Mou et al. 2024), we solely train the temporal blocks, together with the TE and MGF. This strategy seamlessly integrates DiT’s inherent generative knowledge with external motion signals.
@@
  • We introduce Tora, the first trajectory-oriented DiT model for scalable video generation with strong motion guidance. As illustrated in Figure 1, Tora seamlessly integrates various text, visual and trajectory instructions, enabling the creation of motion-manipulable videos.
  • We propose a novel trajectory extractor and a motion-guidance fusion mechanism to facilitate motion control that aligns with the scalability of DiT. Additionally, we ablate several architecture choices and offer empirical baselines for future research.
  • Experiments demonstrate that Tora is capable of generating 720p resolution videos with varying aspect ratios, extending up to 204 frames, all guided by the specified trajectories. Furthermore, it demonstrates superiority in simulating movements within the physical world.
Related Work
Diffusion models for Video Generation
Diffusion models have demonstrated an impressive capability to generate high-quality video samples. Previous research (Ho et al. 2022b, a; Singer et al. 2022; Khachatryan et al. 2023; Zhang et al. 2023b) commonly used video diffusion models (VDMs) that incorporated temporal convolutional and attention layers into the pre-trained image diffusion models. Subsequently, VideoCrafter (Chen et al. 2023) and SVD (Blattmann et al. 2023) expand the application of video diffusion models to larger datasets, while TF-T2V (Wang et al. 2024a) directly learn from extensive text-free videos. Nonetheless, these methods encounter limitations in generating long videos, owing to the inherent constraints on capacity and scalability within the U-Net design. Conversely, DiT-based models (Brooks et al. 2024; Zheng et al. 2024; Bao et al. 2024) can directly generate videos extending up to tens of seconds. Sora (Brooks et al. 2024) converts visual data into a unified representation, facilitating large-scale training and enabling the generation of 1-minute high-definition video. Vidu (Bao et al. 2024) is capable of generating both realistic and imaginative videos in various aspect ratios and resolutions. For our study, we adopt OpenSora (Zheng et al. 2024) as the foundational model, which is an open-source alternative to Sora.
@@
Motion control in Video Generation
To better control motion in generated video, a multitude of studies have endeavored to introduce diverse motion signals in VDMs. Pioneering works like MotionDirector (Zhao et al. 2023) and VMC (Jeong, Park, and Ye 2024) have utilized reference videos to extract motion patterns applicable to diverse video generations. VideoComposer (Wang et al. 2023) expands upon this by adopting depth maps, sketches, or motion vectors from references as conditional inputs for motion control. Nonetheless, these methodologies are limited to reproducing existing motion patterns. Conversely, approaches that leverage trajectories or bounding boxes (Yin et al. 2023; Dai et al. 2023; Wang et al. 2024b) in video generation promise greater adaptability and user accessibility. DragNUWA (Yin et al. 2023) breaks new ground by integrating trajectory-based conditioning into VDMs, facilitating complex camera and object movements. AnimateAnything (Dai et al. 2023) employs motion masks for precise control over the moving regions. TrailBlazer (Wan-Duo Kurt Ma 2023), employs explicit attention mechanisms to maneuver generated objects along precise trajectories. MotionCtrl (Wang et al. 2024b) facilitates more flexible control, allowing separate adjustment of both camera movements and individual object motions. However, all of them yield noticeable artifacts in both motion consistency and visual presentation when applied to longer sequences. In contrast, our method first integrates trajectories into DiT architecture, specifically designed to accommodate scaling properties. This enables closer adherence to the physical world.
Methodology
Preliminary
Latent Video Diffusion Model (LVDM). The LVDM enhances the stable diffusion model (Ramesh et al. 2022) by integrating a 3D U-Net, thereby empowering efficient video data processing. This 3D U-Net design augments each spatial convolution with an additional temporal convolution and follows each spatial attention block with a corresponding temporal attention block. It is optimized employing a noise-prediction objective function:
@@
Here, $\epsilon_{\theta}(\cdot)$ signifies the 3D U-Net’s noise prediction function. The condition $c$ is guided into the U-Net using cross-attention for adjustment. Meanwhile, $z_{t}$ denotes the noisy hidden state, evolving like a Markov chain that progressively adds Gaussian noise to the initial latent state $z_{0}$:
@@
Diffusion Transformer (DiT). The DiT (Peebles and Xie 2023) introduces a novel architecture that merges the strengths of diffusion models with transformer architectures (Vaswani et al. 2017). This integration aims to address the limitations of traditional U-Net-based latent diffusion models (LDMs), improving their performance, versatility, and scalability. While keeping the overall framework consistent with existing LDMs, the key shift lies in replacing the U-Net with a transformer architecture for learning the denoising function $\epsilon_{\theta}(\cdot)$, thereby marking a pivotal advance in the realm of generative modeling.
Tora
@@
During processing, a video autoencoder (Yu et al. 2023a) is first employed to diminish both spatial and temporal dimensions of videos. To elaborate, it encodes the input video $X\in\mathbb{R}^{L\times H\times W\times 3}$ into video latent $z_{0}\in\mathbb{R}^{l\times h\times w\times 4}$, where $L$ denotes the video length and $l=L/4,h=H/8,w=W/8$. $z_{0}$ is next “patchified”, resulting in a sequence of input tokens $I\in\mathbb{R}^{l\times s\times d}$. Here, $s=hw/p^{2}$ and $p$ denotes the patch size. In both SSA and TSA, standard Attention is performed using Query (Q), Key (K), and Value (V) matrices:
$Q=W_{Q}\cdot I_{norm};K=W_{K}\cdot I_{norm};V=W_{V}\cdot I_{norm},$ (3)
Here, $I_{norm}$ is the normalized $I$, $W_{Q},W_{K},W_{V}$ are learnable matrices. The textual prompt is embedded with a T5 encoder and integrated using a cross-attention mechanism.
@@
However, the DiT model employs a video autoencoder and a patchification process to convert the video into patches. Here, each patch is derived across multiple frames, rendering it inappropriate to straightforwardly employ frame-to-frame offsets. To address this, our TE converts the trajectory into motion patches, which inhabit the same latent space as the video patches. Particularly, we begin by transforming the $traj$ into a trajectory map $g\in\mathbb{R}^{L\times H\times W\times 2}$, enhanced with a Gaussian Filter to mitigate scatter. Notably, the first frame employs a fully-zero map. Afterward, the trajectory map $g$ is translated into the RGB color space, producing $g_{vis}\in\mathbb{R}^{L\times H\times W\times 3}$ through a flow visualization technique. We use a 3D VAE to compress trajectory maps, achieving an 8x spatial and 4x temporal reduction, aligning with OpenSora framework. Our 3D VAE is based on the Magvit-v2 architecture, with spatial compression initialized using the VAE of SDXL (Podell et al. 2023) to accelerate convergence. We train the model using only reconstruction loss to obtain the compact motion latent representation $g_{m}\in\mathbb{R}^{l\times h\times w\times 4}$ from the $g_{vis}$.
@@
    Extra channel connections. Denote $h_{i}\in\mathbb{R}^{l\times s\times d}$ as the resultant output from the $i$-$th$ block of the ST-DiT. Following the widespread use of concatenation in GAN-based LVDM, the motion patches are simply concatenated with the previous hidden state $h_{i-1}$ along the channel dimension. An additional convolution layer is then added to maintain the same latent size:
    $h_{i}=Conv~{}([h_{i-1},f_{i}])+h_{i-1},$ (6)

### v4

@@ Abstract
Recent advancements in Diffusion Transformer (DiT) have demonstrated remarkable proficiency in producing high-quality video content. Nonetheless, the potential of transformer-based diffusion models for effectively generating videos with controllable motion remains an area of limited exploration. This paper introduces Tora, the first trajectory-oriented DiT framework that concurrently integrates textual, visual, and trajectory conditions, thereby enabling scalable video generation with effective motion guidance. Specifically, Tora consists of a Trajectory Extractor (TE), a Spatial-Temporal DiT, and a Motion-guidance Fuser (MGF). The TE encodes arbitrary trajectories into hierarchical spacetime motion patches with a 3D motion compression network. The MGF integrates the motion patches into the DiT blocks to generate consistent videos that accurately follow designated trajectories. Our design aligns seamlessly with DiT’s scalability, allowing precise control of video content’s dynamics with diverse durations, aspect ratios, and resolutions. Extensive experiments demonstrate that Tora excels in achieving high motion fidelity compared to the foundational DiT model, while also accurately simulating the complex movements of the physical world. Code is made available at https://github.com/alibaba/Tora .
1 Introduction
Figure 1: More generated samples. Tora effectively manages trajectories to precisely manipulate various objects and backgrounds. In the realm of image-to-video synthesis, it can craft dynamic camera movements in accordance with textual descriptions and the designated starting trajectory points, such as common backgrounds. Furthermore, Tora supports video generation across different aspect ratios, resolutions, and durations, ensuring flexible content creation.
@@ 1 Introduction
Diffusion models [10, 32] have demonstrated their capability to generate diverse and high-quality images or videos. Previously, video diffusion models [14, 4, 51] predominantly employed UNet-based architectures [34], focusing primarily on synthesizing videos of limited duration, typically around two seconds, and were constrained to fixed resolutions and aspect ratios. Recently, Sora [5], a text-to-video generation model leveraging Diffusion Transformer (DiT) [26], has showcased video generation capabilities that significantly outstrip current state-of-the-art methods. Sora excels not only in the production of high-quality videos ranging from 10 to 60 seconds, but also distinguishes itself through its capacity to handle diverse resolutions, various aspect ratios, adherence to the laws of actual physics.
@@
Video generation requires consistent motion across image sequences, underscoring the importance of motion control. Previous works, such as VideoComposer [39] and DragNUWA [48], have implemented generalized motion manipulation through motion vectors and trajectories. Building on this foundation, MotionCtrl [41] innovates by independently managing camera and object motions, thereby expanding the diversity of achievable motion patterns. Despite their promising controllable motion quality, UNet-based methods are restricted to generating videos of only 16 frames at a fixed, lower resolution. This limitation hinders the smooth portrayal of motion, particularly during significant positional shifts in the provided trajectory, leading to distortion and unnatural movements, such as parallel drifting, which diverge from real-world dynamics. Consequently, there is an urgent need for a model capable of producing longer videos with robust motion control and detailed physical representations.
@@
To address these challenges, we present Tora, the first DiT model that simultaneously integrates text, images, and trajectories, enabling scalable video generation with robust motion control. Notably, our work adopts OpenSora [55], an open-source version of Sora, as the foundational DiT model. To align motion control with the scalability of the DiT framework, we propose two novel modules: the Trajectory Extractor (TE), which converts arbitrary trajectories into hierarchical spacetime motion patches, and the Motion-guidance Fuser (MGF), designed to seamlessly integrate these patches within the stacked DiT blocks. More specifically, TE initially converts positional displacements along trajectory into the RGB domain via flow visualization techniques. These visualized displacements undergo Gaussian filtering to mitigate scattered issues. Subsequently, a 3D Variational Autoencoder (VAE) [19] encodes trajectory visualizations into spacetime motion latents, which share the same latent space with video patches. The motion latents are then decomposed into multiple levels of motion conditions via stacked lightweight modules. Our VAE architecture is inspired by MAGVIT-v2 [50] but simplified by omitting codebook dependencies. The MGF integrates adaptive normalization layers [27] to infuse multi-level motion conditions into the corresponding DiT blocks. We explored various adaptations of transformer blocks including adaptive layer normalization, cross-attention, and extra channel connections to inject the motion conditions. Among these, adaptive layer normalization emerged as the most effective to generate consistent videos following the trajectory.
@@
During training, we adapt OpenSora’s workflow to generate high-quality video-text pairs and utilize an optical flow estimator [44] for trajectory extraction. We also integrate a motion segmentor [54] with a camera detector [^1] https://github.com/antiboredom/camera-motion-detector to filter out instances dominated by camera motion, thereby improving our tracking of object trajectories. This careful selection process results in a dataset of high-quality videos with consistent motion. With an adapter-like strategy [24], we solely train the temporal blocks, together with the TE and MGF. This strategy seamlessly integrates DiT’s inherent generative knowledge with external motion signals.
@@
  • We introduce Tora, the first trajectory-oriented DiT model for video generation. As illustrated in Figure 1, Tora enables the creation of motion-manipulable videos with varying aspect ratios, extending up to 204 frames and 720p resolution.
  • We propose a novel trajectory extractor and a motion-guidance fusion mechanism to facilitate motion control that aligns with the scalability of DiT. Additionally, we ablate several architectural choices and scaling capabilities to offer empirical baselines for future research.
  • Experiments demonstrate that Tora achieves state-of-the-art accuracy in controlling object motions. Furthermore, it demonstrates superiority in simulating movements within the physical world.
2 Related Work
2.1 Diffusion models for Video Generation
Diffusion models have demonstrated an impressive capability to generate high-quality video samples. Previous research [14, 13, 36, 17, 52] commonly use video diffusion models (VDMs) that incorporate temporal convolutional and attention layers into the pre-trained image diffusion models. Subsequently, VideoCrafter [7] and SVD [4] expand the application of video diffusion models to larger datasets, while TF-T2V [40] directly learns from extensive text-free videos. Nonetheless, these methods encounter limitations in generating long videos, owing to the inherent constraints on capacity and scalability within the UNet design. Conversely, DiT-based models [5, 55, 3, 47, 29] can directly generate videos extending up to tens of seconds. Sora [5] converts visual data into a unified representation, facilitating large-scale training and enabling the generation of 1-minute high-definition video. Vidu [3] is capable of generating both realistic and imaginative videos in various aspect ratios. CogVideoX [47] introduces an expert diffusion transformer model that generates videos from text prompts or images, along with an effective text-video data processing pipeline that enhances video caption quality.
@@ 2.1 Diffusion models for Video Generation
2.2 Motion control in Video Generation
To better control motion in generated videos, a multitude of studies have endeavored to introduce diverse motion signals in VDMs. Pioneering works like MotionDirector [53] and VMC [16] have utilized reference videos to extract motion patterns applicable to diverse video generations. VideoComposer [39] expands upon this by adopting depth maps, sketches, or motion vectors from references as conditional inputs for motion control. Nonetheless, these methodologies are limited to reproducing existing motion patterns. Conversely, approaches that leverage trajectories [23, 12, 35, 48, 41, 20] or bounding boxes [48, 9, 41] in video generation promise greater adaptability and user accessibility. DragNUWA [48] breaks new ground by integrating trajectory-based conditioning into VDMs, facilitating complex camera and object movements. AnimateAnything [9] employs motion masks for precise control over the moving regions. DragAnything [43] uses the object mask to generate entity representations for achieving motion control. MotionCtrl [41] facilitates more flexible control, allowing separate adjustment of both camera movements and individual object motions. However, all of them yield noticeable artifacts in both motion consistency and visual presentation when applied to longer sequences. In contrast, our method first integrates trajectories into DiT architecture, which enables closer adherence to the physical world.
3 Methodology
3.1 Preliminary
Latent Video Diffusion Model (LVDM). The LVDM enhances the stable diffusion model [32] by integrating a 3D UNet, thereby empowering efficient video data processing. This 3D UNet design augments each spatial convolution with an additional temporal convolution and follows each spatial attention block with a corresponding temporal attention block. It is optimized by employing a noise-prediction objective function:
@@ 3.1 Preliminary
Here, $\epsilon_{\theta}(\cdot)$ signifies the 3D UNet’s noise prediction function. The condition $c$ is guided into the UNet using cross-attention for adjustment. Meanwhile, $z_{t}$ denotes the noisy hidden state, evolving like a Markov chain that progressively adds Gaussian noise to the initial latent state $z_{0}$:
@@
Diffusion Transformer (DiT). The DiT [26] introduces a novel architecture that merges the strengths of diffusion models with transformer architectures [38]. This integration aims to address the limitations of traditional UNet-based latent diffusion models (LDMs), improving their performance, versatility, and scalability. While keeping the overall framework consistent with existing LDMs, the key shift lies in replacing the UNet with a transformer architecture for learning the denoising function $\epsilon_{\theta}(\cdot)$, thereby marking a pivotal advance in the realm of generative modeling.
3.2 Tora
@@
During processing, a video autoencoder [49] is first employed to diminish both spatial and temporal dimensions of videos. To elaborate, it encodes the input video $X\in\mathbb{R}^{L\times H\times W\times 3}$ into video latent $z_{0}\in\mathbb{R}^{l\times h\times w\times 4}$, where $L$ denotes the video length and $l=L/4,h=H/8,w=W/8$. $z_{0}$ is next “patchified", resulting in a sequence of input tokens $I\in\mathbb{R}^{l\times s\times d}$. Here, $s=hw/p^{2}$ and $p$ denotes the patch size. In both SSA and TSA, standard Attention is performed using Query ($Q$), Key ($K$), and Value ($V$) matrices:
$Q=W_{Q}\cdot I_{\mathrm{norm}};K=W_{K}\cdot I_{\mathrm{norm}};V=W_{V}\cdot I_{\mathrm{norm}},$ (3)
Here, $I_{\mathrm{norm}}$ is the normalized $I$, $W_{Q},W_{K},W_{V}$ are learnable matrices. The textual prompt is embedded with a T5 encoder and integrated using a cross-attention mechanism.
@@
However, the DiT model employs a video autoencoder and a patchification process to convert the video into patches. Here, each patch is derived across multiple frames, rendering it inappropriate to straightforwardly employ frame-to-frame offsets. To address this, our TE converts the trajectory into motion patches, which inhabit the same latent space as the video patches. Particularly, we begin by transforming the $traj$ into a trajectory map $g\in\mathbb{R}^{L\times H\times W\times 2}$, enhanced with a Gaussian Filter to mitigate scatter. Notably, the first frame employs a fully-zero map. Afterward, the trajectory map $g$ is translated into the RGB color space, producing $g_{vis}\in\mathbb{R}^{L\times H\times W\times 3}$ through a flow visualization technique. We use a 3D VAE to compress trajectory maps, achieving an 8x spatial and 4x temporal reduction, aligning with OpenSora framework. Our VAE is based on the MAGVIT-v2 architecture, with spatial compression initialized using the VAE of SDXL [28] to accelerate convergence. We train the model using reconstruction loss to obtain the compact motion latent representation $g_{m}\in\mathbb{R}^{l\times h\times w\times 4}$ from the $g_{vis}$.
@@
    Extra channel connections. Denote $h_{i}\in\mathbb{R}^{l\times s\times d}$ as the resultant output from the $i$-$th$ block of the ST-DiT. Following the widespread use of concatenation in GAN-based LVDM, the motion patches are simply concatenated with the previous hidden state $h_{i-1}$ along the channel dimension. An additional MLP is then added to maintain the same latent size:
    $h_{i}=\mathrm{MLP}~{}([h_{i-1},f_{i}])+h_{i-1},$ (6)
@@
3.3 Data Processing and Training Strategy
Data Processing. We employ a structured data processing method to obtain high-quality training videos with consistent object motion. Initially, raw videos are segmented into shorter clips based on scene detection [^2] https://github.com/Breakthrough/PySceneDetect. Subsequently, we remove invalid videos, such as those with encoding errors, zero duration, or low resolution. Furthermore, we utilize aesthetic [^3] https://github.com/christophschuhmann/improved-aesthetic-predictor and optical flow scores [44] to filter out low-quality videos. To concentrate on the motion of primary objects, we implement camera motion filtering, using results from motion segmentation [54] and camera detection to exclude instances predominantly exhibiting camera movement. Dramatic object motions in certain videos can lead to significant optical flow deviations, which may interfere with trajectory training. To address this, we retain these videos based on a probability of $(1-flow\_score/100)$. For eligible videos, we generate captions using the PLLaVA model [45]. During inference, we utilize GPT-4o [25] to refine prompts, ensuring alignment with training process.
Motion condition training. Inspired by DragNUWA and MotionCtrl, we adopt a two-stage training approach for trajectory learning. In the first stage, we extract dense optical flow [44] from the training video as the trajectory, providing richer information to enhance motion learning. In the second stage, we adjust the model from complete optical flow to more user-friendly trajectories by randomly selecting 1 to $N$ object trajectories based on motion segmentation results and flow scores. To improve the scattered nature of sparse trajectories, we apply a Gaussian filter for refinement. After completing the two-stage training, Tora facilitates flexible motion control using arbitrary trajectories.
Image condition training. We employ a mask strategy to support visual conditioning. Specifically, we randomly unmask frames during training, ensuring that the video patches of unmasked frames are not subjected to any noise.
4 Experiments
4.1 Experimental Setup
Implementation Details. Tora is initialized with OpenSora v1.2 weights, and training videos have resolutions from 144p to 720p and frame counts ranging from 51 to 204. To balance training FLOP and memory usage, we adjust the batch size from 1 to 50. We use Adam Optimizer [18] with a learning rate of $2\times 10^{-5}$ on 8 NVIDIA A100. The 3D VAE is initially trained on datasets  [22, 21, 33, 6] for optical flow estimation and then frozen during Tora training. We train Tora for 2 epochs with dense optical flow and fine-tune for 1 epoch with sparse trajectories. The maximum number of sampling trajectories $N$ is set to 16. The inference step and the guidance scale are set to 30 and 7.0, respectively.
Dataset. Our training videos are sourced from four datasets: 1) Panda-70M [8], from which we use the training-10M subset containing high-quality videos; 2) Mixkit [11]; 3) Pexels [2]; and 4) Internal videos. The internal videos are manually annotated, with each clip labeled to include object masks and camera movement. Following our data processing pipeline, we select about 630k eligible videos for the training dataset. For inference, we curate 185 clips with diverse motion trajectories and scenes, to serve as a new benchmark for evaluating the motion controllability.
Metrics. We leverage standard metrics including Fréchet Video Distance (FVD) [37], and CLIP Similarity (CLIPSIM) [42] to quantitatively evaluate video quality. For assessing motion controllability, we utilize the Trajectory Error (TrajError) metric, which calculates the average L1 distance between the generated and predefined trajectories.
4.2 Results